# Membership Privacy Risks of Sharpness Aware Minimization

### v3

@@ 4.2 SAM Increases Influence of High Memorization Samples
Figure 2: (a) and (b): Distribution of the influence scores of the 20 training samples achieving the highest influence score over each memorization interval over their memorization scores for SGD (a) and SAM (b). The regression analysis (green lines) shows that SAM maintains a smoother influence distribution, amplifying mid-to-high memorization samples (subclass features), while SGD relies more heavily on a narrow set of highly memorized points (noise). (c) Difference in influence scores between SAM and SGD as a function of memorization score differences. SAM downweights low-memorized samples and selectively amplifies the influence of mid-to-high memorization samples.
We here analyze how memorization affects the influence of training samples on test predictions, using the influence metric defined in Equation˜2 for both SGD and SAM. Following the setup of Feldman & Zhang (2020), we first filter training–test sample pairs with influence scores above 0.2 to exclude non-influential cases. We then group training samples by memorization intervals – defined as $l<\text{mem}(\mathcal{A},\mathcal{D},i)<u$, with $l$ and $u$ ranging from 0 to 1 in steps of 0.05 — and, for each interval, we select the 20 training samples achieving the highest influence score on test data. This yields a distribution of influence scores of the most influential training samples conditioned on their memorization levels.
Figures˜2(a) and 2(b) show the resulting distributions for SGD and SAM, respectively. As in Figure˜1(b), we fit a regression curve (green line) to highlight the trends. For SGD, influence scores incur in a steep transition from lowly influent samples to highly influent points at the upper end of the memorization spectrum. This indicates SGD’s reliance on a very narrow set of highly memorized samples. In contrast, SAM exhibits a smoother transition curve, with a larger set of high (and mid-to-high) memorization samples contributing more consistently to test predictions. This supports our earlier finding that SAM emphasizes a set of atypical, non-dominant subclass patterns rather than focusing on sample-specific noise.
@@ 4.3 SAM’s Generalization Gain Comes From Higher Memorization of Sub-patterns
We measure the typicality of a test data point as the entropy of its corresponding training samples’ influence scores. We rely on this measure since, in practice, the prediction of a typical unseen sample would be evenly influenced by numerous training data points within the same class (for both SGD and SAM), while atypical counterparts would be heavily influenced by a handful of training samples that are themselves also atypical (for SGD) or by a larger set of training samples containing atypical sub-patterns (for SAM). To measure the even spread, we normalize the influence scores and leverage entropy. Even influence spread would follow a more uniform distribution resulting in high entropy, while uneven spread incurs low entropy. Formally, for each test data point $i$, let $\mathcal{S}_{i}$ be the set of influence scores of all the training points in the same class. Let $S_{i,j}$ be the influence score of $j_{th}$ training point and $m$ be the number of training points in the same class. Then, our entropy metric $\mathcal{I}_{ent}$ is defined as:
$\mathcal{I}_{ent}[i]=\sum_{j=1}^{m}-p_{i,j}*\log{p_{i,j}},\text{ where, }p_{i,j}=\frac{\mathcal{S}_{i,j}}{\sum_{j}\mathcal{S}_{i,j}}$ (6)
We group test data points into 5 buckets in the order of lowest $\mathcal{I}_{ent}$ to highest $\mathcal{I}_{ent}$. We present some test images and their top-10 influential training images in Figure˜3 from bucket 1 and bucket 5. The figure illustrates that images from bucket 1 tend to be atypical images – e.g., bicycle alongside people, white tiger, etc, – for their respective labels while images from bucket 5 tend to be more typical images—e.g., typical bicycle and yellow tiger. For quantitative verification, we plot the distribution of memorization scores of the highest influencing training points from each bucket. We observe that lower (higher) numbered buckets are influenced by training points with higher (lower) memorization scores (see Figures˜4(b) and 4(c)). The results for other buckets interpolate between those of bucket 1 and 5, and are skipped for brevity.
We compare the generalization gains of SAM against SGD on each of these buckets and show the results in Figure˜4(a) [^3] These results do not consider image transformations (e.g. random crop, rotations), however we have also replicated the experiment with transformations, obtaining a similar trend.. For test data points in bucket 5, SGD achieves a negligible performance gain, while for bucket 1 SAM achieves a significant gain w.r.t. SGD. Thus, the performance gains of SAM can be attributed to it correctly predicting more atypical data points which need more memorization of atypical sub-patterns to be classified correctly. For further validation, we generate a synthetic dataset and illustrate SAM’s capability of learning atypical subclasses better than SGD in Appendix˜B.
@@
These results – together with the ones obtained in Sections˜4.1 and 4.2 – provide strong empirical evidence that SAM’s increased performance derives from an increased targeted memorization which focuses on atypical but informative subclass patterns that enhance generalization over atypical test data. This explains both SAM’s improved test performance and its increased vulnerability to membership inference attacks, as these memorized patterns make training samples more distinguishable from unseen data.
@@ 5 Theoretical Analysis
In this section, we theoretically analyze a setting involving majority subclass samples, minority subclass samples, and pure noise samples to illustrate how an overfitting model that generalizes better can also carry higher membership privacy risk. Driven by the motivations in the previous sections, we discern a model by the amount of minority subclass feature it captures. Theorem˜1 shows that this property leads to higher generalization. Moreover, Theorem˜2 shows that this property can also lead to higher membership privacy risk, completing our analysis. We consider perfectly interpolating linear predictors for a binary classification task. The full proofs are deferred to Appendix˜C.
@@
Now, we formally define attack setup and attacker’s advantage for MIA. We focus on the confidence threshold based attack as it is empirically one of the most effective attacks and naturally connected to entropy and modified entropy.
Definition 3 (Confidence-threshold attacker).
For $t>0$ (probabilistic confidence $\tau=\sigma(t)$), define $\mathcal{A}_{t}(x,y)=\mathds{1}\{y\langle\mathbf{w},\mathbf{x}\rangle\geq t\}$ and the attacker’s advantage
$\mathrm{Adv}^{\mathrm{M}}(w,t)\;\coloneqq\;\Pr\big(\mathcal{A}_{t}{=}1\mid(\mathbf{x},y)\in D\big)\;-\;\Pr\big(\mathcal{A}_{t}{=}1\mid(\mathbf{x},y)\sim\mathcal{D}\big).$
where the probability is taken over S.
Note that the attacker’s advantage is the difference in the True Positive Rate (TPR) and the False Positive Rate (FPR). Now we state our main results.
@@
Theorem˜1 shows that the interpolating model that aligns with the minority subclass feature more generalizes better.
Theorem 2 (Higher MSA $\implies$ Higher Attacker’s Advantage).
Let $\mathbf{w}^{(A)},\mathbf{w}^{(B)}$ be two interpolating solutions trained on the same $D$ and $\mathbf{w}^{(A)}\;\overset{\mathrm{MSA}}{\succcurlyeq}\;\mathbf{w}^{(B)}$. Choose any threshold $t=$ logit($\tau$) such that
$\max\{A^{(A)},A^{(B)},m_{0}^{(A)},m_{0}^{(B)}\}\;<\;t$
Then under Definition˜1 and regulatory conditions,
$\mathrm{Adv}^{\mathrm{M}}\!\big(\mathbf{w}^{(A)},t\big)\;\geq\;\mathrm{Adv}^{\mathrm{M}}\!\big(\mathbf{w}^{(B)},t\big),$
with strict inequality whenever some minority training indicator at threshold $t$ flips under $\mathbf{w}^{(A)}$ but not under $\mathbf{w}^{(B)}$.
Theorem˜2 shows that the model that exhibits higher generalization under our setting can also exhibit higher attacker’s advantage. The condition on the threshold $t$ is easily met in practice as the learned probability threshold $\sigma(t)$ is often high (between 0.9 and 0.98 in our experiments).
6 Conclusion and Future Work
Our work is an exploration of SAM’s properties pertaining to long-tailed data distributions to unearth surprising tradeoffs between membership privacy risk and generalization. With the insights offered by our experiments and theoretical analysis, we hope to motivate more privacy aware methods in the future.
Appendix A LLM usage
LLMs were used to assist with grammar correction and sentence-level proofreading throughout the manuscript.
Appendix B Synthetic Dataset
(a) Class contours
(b) SAM vs SGD (Few atypical examples)
(c) SAM vs SGD (Large sample size)

### v4

@@ 4.2 SAM Increases Influence of High Memorization Samples
Figure 2: (a) and (b): Distribution of the influence scores of the 20 most influential training samples over each memorization interval for SGD (a) and SAM (b). The regression analysis (green lines) shows that SAM maintains a smoother influence distribution, relying more on mid-to-high (0.6 - 0.85) memorization samples (subclass features) than SGD. (c) Difference in influence scores between SAM and SGD as a function of memorization score differences. SAM downweighs influence for low-memorized samples and selectively amplifies the influence of mid-to-high memorization samples.
We here analyze how memorization affects the influence of training samples on test predictions, using the influence metric defined in Equation˜2. Following the setup of Feldman & Zhang (2020), we first filter training–test sample pairs with influence scores above 0.2 to exclude non-influential cases. We then group training samples by memorization intervals – defined as $l<\text{mem}(\mathcal{A},\mathcal{D},i)<u$, with $l$ and $u$ ranging from 0 to 1 in steps of 0.05 — and, for each interval, we select the 20 training samples achieving the highest influence score on test data. This yields a distribution of influence scores of the most influential training samples conditioned on their memorization levels.
Figures˜2(a) and 2(b) show the resulting distributions for SGD and SAM, respectively. As in Figure˜1(b), we fit a regression curve (green line) to highlight the trends. For SGD, influence scores incur in a steep transition from lowly influent samples to highly influent points at the upper end of the memorization spectrum. This indicates SGD’s reliance on a very narrow set of highly memorized samples. In contrast, SAM exhibits a smoother transition curve, with a larger set of high (and mid-to-high) memorization samples contributing more consistently to test predictions. This supports our earlier finding that SAM emphasizes a set of atypical, non-dominant subclass patterns.
@@ 4.3 SAM’s Generalization Gain Comes From Higher Memorization of Sub-patterns
We measure the typicality of a test data point as the entropy of its corresponding training samples’ influence scores. We rely on this measure since, in practice, the prediction of a typical unseen sample would be evenly influenced by numerous training data points within the same class, while atypical counterparts would be heavily influenced by a handful of training samples that are themselves also atypical. To measure the even spread, we normalize the influence scores and leverage entropy. Even influence spread would follow a more uniform distribution resulting in high entropy, while uneven spread incurs low entropy. Formally, for each test data point $i$, let $\mathcal{S}_{i}$ be the set of influence scores of all the training points in the same class. Let $S_{i,j}$ be the influence score of $j_{th}$ training point and $m$ be the number of training points in the same class. Then, our entropy metric $\mathcal{I}_{ent}$ is defined as:
$\mathcal{I}_{ent}[i]=\sum_{j=1}^{m}-p_{i,j}\cdot\log{p_{i,j}},\text{ where, }p_{i,j}=\frac{\mathcal{S}_{i,j}}{\sum_{j}\mathcal{S}_{i,j}}$ (5)
We group test data points into 5 buckets in the order of lowest $\mathcal{I}_{ent}$ to highest $\mathcal{I}_{ent}$. We present some test images and their top-10 influential training images in Figure˜3 from bucket 1 and bucket 5. The figure illustrates that images from bucket 1 tend to be atypical images – e.g., bicycle alongside people, white tiger, etc, – for their respective labels while images from bucket 5 tend to be more typical images—e.g., typical bicycle and yellow tiger. For quantitative verification, we plot the distribution of memorization scores of the highest influencing training points from each bucket. We observe that lower numbered buckets are associated with high memorization and vice versa (see Figures˜4(b) and 4(c)). The results for other buckets interpolate between those of bucket 1 and 5, and are skipped for brevity.
We compare the generalization gains of SAM against SGD on each of these buckets and show the results in Figure˜4(a) [^4] These results do not consider image transformations (e.g. random crop, rotations), however we have also replicated the experiment with transformations, obtaining a similar trend.. For test data points in bucket 5, SGD achieves a negligible performance gain, while for bucket 1 SAM achieves a significant gain w.r.t. SGD. Thus, the performance gains of SAM can be attributed to it correctly predicting more atypical data points which need more memorization of atypical sub-patterns to be classified correctly. For further validation, we generate a synthetic dataset and illustrate SAM’s capability of learning atypical subclasses better than SGD in Appendix˜A. We theoretically analyze how stronger alignment with minority subclass features can lead to better generalization in Section B.
@@
These results – together with those in Sections˜4.1 and 4.2 – provide strong empirical evidence that SAM’s increased performance derives from better capturing atypical but informative sub-patterns, allowing predictions to be less dominated by the majority feature. For SGD, we posit that this results in a higher output sensitivity, yielding more confident predictions on unseen data. Paradoxically, this high variance acts as a cloak for membership privacy, as it mimics the high confidence usually reserved for members. In the next section, we theoretically analyze how sharpness-regularization can inherently suppress prediction variance on unseen data.
@@ 5 Theoretical Analysis
In this section, we provide a theoretical foundation for the variance–shrinkage effect of sharpness-aware minimization and its implication for membership inference risk. To obtain a clean and analyzable characterization of SAM’s geometric bias, we study perfectly interpolating regime of overparameterized linear models. Closest to our analysis in spirit is that of Tan et al. (2022), which explains how parameter size and ridge regression affects membership risk via a variance gap. In contrast, we introduce a curvature-aligned geometry modeling SAM and show that it provably increases membership risk.
Our proofs are written in the regression form $X\theta=y$. We note that the connection to classification is direct. For separable binary classification with losses such as logistic or exponential, gradient descent is known to implicitly maximize the $\ell_{2}$ margin, and the resulting predictor is geometrically equivalent to the minimum-$\ell_{2}$-norm interpolator of a corresponding regression problem (Soudry et al., 2024; Gunasekar et al., 2017; Muthukumar et al., 2021). Motivated by this equivalence, we can consider classification as regression to high-magnitude targets $y_{i}\in\{-M,+M\}$ for a large constant $M\gg 1$ under MSE loss. In this regime, training points are interpolated to $\pm M$ and therefore have fixed high confidence, whereas test points yield outputs that fluctuate around the decision boundary (0).
Model and notation.
We work in finite dimension $d\gg n$. Let the population feature covariance be $\Sigma\in\mathbb{R}^{d\times d}$ with $\Sigma\succ 0$. Draw i.i.d. samples
$x_{i}\sim\mathcal{N}(0,\Sigma),\qquad y_{i}=\theta^{*\top}x_{i}+\xi_{i},\quad\xi_{i}\sim\mathcal{N}(0,\sigma_{y}^{2}),\ \text{independent of }x_{i},\quad i=1,\dots,n.$
Let $X\in\mathbb{R}^{n\times d}$ have rows $x_{i}^{\top}$, $y=(y_{1},\dots,y_{n})^{\top}$, and assume $\mathrm{rank}(X)=n$. Define the orthogonal projector onto the data span and the covariance matrix
$P:=X^{\top}(XX^{\top})^{-1}X,\qquad P^{2}=P=P^{\top},\qquad\widehat{\Sigma}:=\frac{1}{n}X^{\top}X,\qquad\Sigma:=\mathbb{E}[xx^{\top}].$
The model is defined as $f_{G}(x):=\widehat{\theta}_{G}^{\top}x$. In the squared-loss linear model, the population Hessian $H$ equals $\Sigma$ and the empirical Hessian $\widehat{H}$ equals $\widehat{\Sigma}$.
Now we introduce the geometries we employ to model SGD and SAM.
Min-$G$ interpolation.
For any symmetric positive definite matrix $G\succ 0$, consider the minimum-$G$-norm interpolant:
$\widehat{\theta}_{G}\ :=\ \arg\min_{\theta\in\mathbb{R}^{d}}\ \frac{1}{2}\,\theta^{\top}G\,\theta\quad\text{s.t.}\quad X\theta=y.$ (6)
We compare the standard Euclidean case,
$G_{0}:=I_{d}$
with the sharpness-aware geometry,
$G_{\eta}:=I+\eta H=I+\eta\Sigma,\qquad\eta>0,$
Classical implicit-bias results identify standard SGD with the Euclidean case under step size and weight initialization assumptions (Zhang et al., 2017). Sharpness-aware geometry, on the other hand, reflects SAM’s penalty on local sharpness, which can be shown by expanding the minimization objective. We prove that, relative to the Euclidean interpolation, the geometry $G_{\eta}$ strictly reduces the variance of non-member outputs, thereby enlarging the separation between members and non-members and increasing the advantage of membership inference attacks.
We first establish that SAM strictly reduces the variance of the model’s output on unseen data. By Lemma 3, for a non-member sample $X_{\mathrm{out}}$, and the output follows $f_{G}(X_{\mathrm{out}})\ \sim\ \mathcal{N}\!\big(0,\ \sigma_{G}^{2}\big)$.
Theorem 1 (Variance strictly decreases for SAM geometry).
Under Assumption 1, there exists $\eta_{0}>0$ such that for all $\eta\in(0,\eta_{0}]$,
$\sigma_{G_{\eta}}^{2}<\sigma_{G_{0}}^{2}\qquad\text{with probability }1-o(1).$
Remark 1.
Theorem 1 formalizes a simple geometric picture. The minimum-$G$ interpolant balances fitting the training span against penalizing components in directions where $G$ is large. Moving from $G_{0}$ to $G_{\eta}$ increases the penalty precisely along high-curvature directions of the Hessian. Lemma 2 shows that, under the overlap condition in Assumption 1, this reweighting strictly suppresses the $(I-P)\Sigma\widehat{\theta}_{G_{0}}$ component of the interpolant. Since non-member predictions depend on $\widehat{\theta}_{G}$ only through the quadratic form $\widehat{\theta}_{G}^{\top}\Sigma\widehat{\theta}_{G}$, this suppression translates directly into a strict decrease of the output variance on unseen data.
Next, we consider confidence-threshold attack and likelihood ratio attack.
Confidence-threshold attack.
The attacker uses the confidence score $\mathrm{Conf}_{G}(x):=|f_{G}(x)|$ and predicts “member” iff $\mathrm{Conf}_{G}(x)\geq\tau$. Let $(X_{\mathrm{in}},Y_{\mathrm{in}})$ be a random training pair and $X_{\mathrm{out}}\sim\mathcal{N}(0,\Sigma)$ an independent non-member. Define attack advantage
$\mathrm{Adv}^{\mathrm{conf}}_{G}:=\sup_{\tau\geq 0}\big(\mathrm{TPR}_{G}(\tau)-\mathrm{FPR}_{G}(\tau)\big).$
Theorem 2 (SAM strictly increases confidence-based MIA advantage).
Under Assumption 1, for all sufficiently small $\eta>0$,
$\mathrm{Adv}^{\mathrm{conf}}_{G_{\eta}}\;>\;\mathrm{Adv}^{\mathrm{conf}}_{G_{0}}\qquad\text{with probability }1-o(1).$
Remark 2.
For the confidence-threshold attack, interpolation implies that member confidences are geometry-invariant: the training logits are fixed (up to label noise) for any choice of $G$. Thus, changing the geometry from $G_{0}$ to $G_{\eta}$ leaves $\mathrm{TPR}_{G}(\tau)$ unchanged for every threshold $\tau$, while Theorem 1 strictly reduces the non-member variance and hence lowers $\mathrm{FPR}_{G}(\tau)$ at every $\tau>0$. In other words, SAM pushes non-member scores closer to the decision boundary, sharpening the separation between the two confidence distributions.
Likelihood Ratio (LR) attack.
The oracle LRA predicts “member” iff $\Lambda_{G}(f_{G}(x))\geq t$. Define
$\Lambda_{G}(s):=\log\frac{p_{\mathrm{in}}(s)}{p_{\mathrm{out}}(s;G)},\qquad\mathrm{Adv}^{\mathrm{LR}}_{G}:=\sup_{t\in\mathbb{R}}\big(\Pr(\Lambda_{G}(f_{G}(X_{\mathrm{in}}))\geq t)-\Pr(\Lambda_{G}(f_{G}(X_{\mathrm{out}}))\geq t)\big).$
Theorem 3 (SAM strictly increases LR-attack advantage).
Under Assumption 1, for all sufficiently small $\eta>0$,
$\mathrm{Adv}^{\mathrm{LR}}_{G_{\eta}}\;>\;\mathrm{Adv}^{\mathrm{LR}}_{G_{0}}\qquad\text{with probability }1-o(1),$
where $G_{\eta}=I+\eta\Sigma$.
Remark 3.
The argument is similar to Theorem 2. Theorem 3 analyzes an oracle LR attacker that uses a single global IN/OUT distribution. Corollary 5 strengthens this to sample-adaptive IN/OUT distributions, matching the per-query calibration used by LiRA/RMIA. We note that while Theorem 2 and Theorem 3 state an improvement in the supremum advantage, the proof in fact shows a stronger statement: for all sufficiently small $\eta>0$, the entire ROC curve of the confidence-based attack under $G_{\eta}$ strictly dominates that under $G_{0}$. This is empirically verified in Figure 7, where SAM’s ROC curve is above that of SGD’s for nearly the entire range across most settings.
6 Conclusion and Future Work
This work investigates the mechanism behind SAM’s dual effect: superior generalization coupled with heightened privacy leakage. As algorithms seeking flatter minima are widely employed, our work serves as a cautionary tale users should be aware of. Furthermore, extending our findings to devise an effective defense against Membership Inference Attacks would be beneficial to the community. Finally, future research is encouraged to further scrutinize the implicit bias of SAM.
Appendix A Synthetic Dataset
(a) Class contours
(b) SAM vs SGD (Few atypical examples)
(c) SAM vs SGD (Large sample size)
Appendix B Better minority subclass alignment leads to higher generalization
@@
In this section, we theoretically analyze a setting involving majority subclass samples, minority subclass samples, and pure noise samples to illustrate how an overfitting model generalizes better. Driven by the motivations in the previous sections, we discern a model by the amount of minority subclass feature it captures. Theorem˜4 shows that this property leads to higher generalization. The proofs are in Section C.3.
@@
Theorem˜4 shows that the interpolating model that aligns with the minority subclass feature more generalizes better.